The velocity of an enzyme is 20% Vmax. What is the ratio of [S] to KM...

  1. The velocity of an enzyme is 20% Vmax. What is the ratio of [S] to KM under this condition?
Answer #1

since v is 20% of Vmax assume Vmax=1, therefore v= 20/100 * 1 = 0.20

We have the relation

1/v = Km/Vmax x 1/[S] + 1/Vmax
1/0.20 = Km/1 x 1/[S] + 1/1
5 - 1 = Km/[S]
4 = Km/[S]

1/4 = 0.25 = [S] / Km